OpenStudy (anonymous):

If 2 < 20x - 13 <3, what is one possible value for x?

OpenStudy (anonymous):

Answer I get after solving 0.75 < x < 0.8 , so I need to know the one posible value for x now?

OpenStudy (anonymous):

2+13<20x<3+13 15<20x<16 15\20<x<16\20 3\4<x<4\5

OpenStudy (anonymous):

Yes, so now I need to know the one posible value for x now?

OpenStudy (anonymous):

(P.s. Oh yes I knew how to work it out, I just dont get what os the one possible value for x?)

jimthompson5910 (jim_thompson5910):

You can pick any number that lies between 3/4 and 4/5. So why not the midpoint (3/4+4/5)/2 = (15/20 + 16/20)/2 = (31/20)/2 = 31/40 This means that the number 31/40 is the midpoint between 3/4 and 4/5 So one possible value for x is x = 31/40

OpenStudy (anonymous):

The answer at the back of he book is 76 for this question. I think something is wrong.

OpenStudy (anonymous):

Either in the question, or the answer.

jimthompson5910 (jim_thompson5910):

Well 76 is clearly outside the bounds, so yes that's my guess too.

OpenStudy (anonymous):

My guess is the book says .76, either you misread the decimal point or the book has typo :)

OpenStudy (anonymous):

Okay, even if I take 0.76 as the value for x? I would ask WHY?

OpenStudy (anonymous):

Why 0.76?

jimthompson5910 (jim_thompson5910):

Because 0.76 is larger than 0.75 and smaller than 0.8 So 0.75 < 0.76 < 0.8

jimthompson5910 (jim_thompson5910):

And 0.76 = 0.75 + 0.01, so that may be why the book chose this number

OpenStudy (anonymous):

Obviously .76 belongs to the range [ .75, .8 ]

OpenStudy (anonymous):

okay thanks.